OpenStudy (anonymous):

4/3(c+4) + 1 = 2c/(c+4)

OpenStudy (anonymous):

c = (-29-5 i sqrt(15))/8 or c = (-29+5 i sqrt(15))/8

OpenStudy (gw2011):

4/3(c+4) + 1 = 2c/(c+4) 4/3(c+4) - 2c/(c+4) + 1 = 0 [4 - 2c(3) +3(c+4)]/3(c+4) = 0 (4 - 6c +3c +12)/3(c+4) = 0 (-3c+16)/3(c+4) = 0 Multiply both sides by 3(c+4) to eliminate denominator -3c+16 = 0 -3c = -16 c = 16/3 To verify if c=16/3 is correct substitute 16/3 for c for both the left side and right side of the equation: 4/3(c+4) + 1 4[(3)(16/3+4)+1=4/(16+12) + 1=4/28+1=1/7+1=8/7 2c/(c+4) (2)(16/3)/(16/3+4)=(32/3)/(28/3)=(32/3)(3/28)=8/7